OpenStudy (anonymous):

Solve the logarithmic equation algebraically. log(4x) – log(13 + √x) = 2

OpenStudy (anonymous):

use log a - log b = log(a/b) and 2 = log 100

OpenStudy (anonymous):

Ok so using the rule: log(x) - log(y) = log(x/y) We can establish this from your equation.